# Study of symmetries in finite temperature $N_{f}=2$ QCD with Möbius Domain Wall Fermions

### v1

@@ 1 Introduction
Study of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ (broken spontaneously by quantum anomaly) in $N_{f}=2$ is an important supplement to the study of Quantum Chromodynamics (QCD) with nondegenerate flavor such as the studies done by [1, 2, 3, 4, 5, 6]. The core physics of $SU(N_{f})_{L}\times SU(N_{f})$ primarily relies on $N_{f}=2$ as all nondegenerate flavors have large masses and will explicitly break the $SU(N_{f})$ symmetry down to the two lowest mass quarks. Therefore, direct study of the chiral phase transition in 2-Flavor QCD is quite straightforward with a core physical signature which is replicated in QCD with larger flavor symmetry. Chiral symmetry breaking is of interest to ongoing work concerning the phases of QCD, as it is broken spontaneously at low temperatures and may be related to higher temperature symmetries associated with deconfinement through the chiral condensate. It has also been argued that the chiral condensate is not simply an order parameter for chiral symmetry transitions, but may in fact have a relationship to the susceptibility to the axial anomaly $U(1)_{A}$ which is broken anomalously below the scale of QCD. While the anomaly has been considered a low energy effect, there is some evidence[7, 8, 9, 10, 11, 12] that the chiral condensate is connected to reduced susceptibility to the quantum anomaly at high temperatures. The relationship of these two symmetries, and their transitions around the critical point are important to the continued study of high temperature QCD and deconfined matter which were present at early times in the universe.
@@
2 Möbius Domain Wall Fermions
The four dimensional fermion action is described by an approximation of the sign function in the four dimensional overlap operator, we will use the form of Nueberger’s overlap operator:
$D_{OV}=\frac{1+m}{2}-\frac{1-m}{2}\gamma_{5}\mbox{sgn}(H).$ (1)
Our choice of Möbius domain wall fermions is related to our choice in the form of the kernel function $H$ which is
$H=\frac{\gamma_{5}D_{W}}{2+\gamma_{5}D_{W}}$
For a sufficiently large but finite fifth dimensional extent $L_{s}$ we can approximate the function $\mbox{sgn}(H)$ as $\tanh(H)$ like under the assumption that the eigenvalue spectrum of the kernel $H$ is of $\mathcal{O}(1)$.
$\epsilon(H)=\dfrac{(H+1)^{L_{s}}-(H-1)^{L_{s}}}{(H+1)^{L_{s}}+(H-1)^{L_{s}}}=\tanh(L_{s}\tanh^{-1}(H)).$ (2)
For our study we take $L_{s}=16$ which reduces the residual mass or violation of the Ginsparg-Wilson relation to $<1~{}\rm{MeV}$. This minimizes the symmetry breaking effects for screening masses for $\mathcal{O}(100~{}\rm{MeV})$ and above, assuring that all of our meson correlators are chiral at the temperatures in our range of study. The key benefit for this work is that we have a theoretically clean approach to analysis of correlation functions because symmetry transformations applied to each channel can be studied from the data directly [27, 28, 29].
@@ 3.1 Spatial Mesonic Correlators
As the quarks are subject to anti-periodic boundary conditions in the temporal direction, the momentum recieves a large contribution from the Matsubara frequencies. If the largest contribution to the two-point correlation function is from a pair of static non-interacting quarks, then the screening mass $M_{\Gamma}$ is well approxmiated to $2\pi T$, twice the lowest Matsubara frequency. Therefore, a comparison between these two quanitites $M_{\Gamma}$ and $2\pi T$ provides an effective estimation of the strength of interaction between quarks in QCD.
@@ 3.2 Emergence of $SU(2)_{CS}$
The set of the generators of the so-called $SU(2)_{CS}$ chiral-spin or $SU(2)_{CS}$ symmetry, explored in [17, 16, 18, 30, 31]. However, for the case of QCD with quarks interacting with the gluonic medium the symmetry does not appear as directly as the free quark example above, and appears to be approximate. This is due to contributions from the chromoelectic term of the quark-gluon interaction(when fixed to the rest frame of the medium), as this term is symmetric under the same transformations $\Sigma$. However, the symmetry breaks down at higher temperatures due the increase in chromomagnetic effects and thus the emergent symmetry of the chromoelectic and quarks field breaks in the high temperature limit[30, 1, 2, 5].
@@ 4 Numerical Results
Previous work on $N_{f}=2$ QCD have also considered an $C(z)\sim\exp(-mz)/z$ fitting form, which is predicted from the free two quark propagator; however, for the low temperatures in this study, the free quark fitting ansatz cannot be applied.
Figure 1 is an example of effective mass for the $PS$ 1(a) and $T$ 1(b) channels, which are taken from our lightest mass $m=0.0010$ ensemble at temperature $T=147~{}\rm{MeV}$. For fittings which have significant noise, both the effective mass fits as well as correlator fits were used to determine a reliable fitting range.
Figure 1: Above are the effective mass plots for the $PS$ and $Tt$ channels with $m=0.0010$ at $T=147~{}\rm{MeV}$.
@@
Figure 7: Multiple quark masses are plotted for the highest temperature lattice $T=330~{}\rm{MeV}$, for the all quarks $SU(2)_{CS}$ remains broken.
5 Summary of Preliminary Results
In figure 2 the fitted mass versus temperature is plotted for the lightest mass quarks. Due to the difference in the psuedocritical transition temperature for $N_{f}=2$ and $N_{f}=2+1$ QCD figure 3 is the plot of temperature normalized by $T_{c}$ and fitted masses normalized by $2\pi T$. This allows a direct comparison between the 2019 results from HotQCD in [32] and our 2-flavor study.
The most significant difference between our work and the results by HotQCD, which are plotted as shaded bands in figs 2 and 3, is the behavior of the $S$ channel at low temperatures. Because our simulations are done with Möbius Domain Wall fermions we do not observe the same $2\pi^{\pm}$ lattice artifact, and instead, we observe a large increase in the screening mass in the region of broken chiral symmetry.
In contrast to both of these symmetries we also plotted a temperature spectrum for the $SU(2)_{CS}$ predicted in [15, 16, 17], unlike the clear signals of symmetry restoration seen in figs 4 and 5, the plot of the symmetry pair $X-A$ figure 5 does not converge to zero for the temperatures in our study. While there is a significant reduction in noise consistent with the symmetry restoration for the other symmetry pairs, we do not see a convergence to, or sustained, zero value indicating emergence of $SU(2)_{CS}$. This is seen also in figure 7 which shows $SU(2)_{CS}$ in comparison to other symmetries at $T=330~{}\rm{MeV}$ for which there is not even an approximate threshold as the channels connected by $SU(2)_{CS}$ remain significantly broken, which is true for all temperatures in this study.

### v2

@@ 1 Introduction
Study of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ (broken spontaneously by quantum anomaly) in $N_{f}=2$ is an important supplement to the study of Quantum Chromodynamics (QCD) with nondegenerate flavor such as the studies done by [1, 2, 3, 4, 5, 6]. The core physics of $SU(N_{f})_{L}\times SU(N_{f})$ primarily relies on $N_{f}=2$ as all nondegenerate flavors have large masses and will explicitly break the $SU(N_{f})$ symmetry down to the two lowest mass quarks. Therefore, direct study of the chiral phase transition in 2-Flavor QCD is quite straightforward with a core physical signature which is replicated in QCD with larger flavor symmetry. Chiral symmetry breaking is of interest to ongoing work concerning the phases of QCD, as it is broken spontaneously at low temperatures and may be related to higher temperature symmetries associated with deconfinement through the chiral condensate. It has also been argued that the chiral condensate is not simply an order parameter for chiral symmetry transitions, but may in fact have a relationship to the susceptibility to the axial anomaly $U(1)_{A}$ which is broken anomalously below the scale of QCD. While the anomaly has been considered a low energy effect, there is some evidence[7, 8, 9, 10, 11, 12, 13] that the chiral condensate is connected to reduced susceptibility to the quantum anomaly at high temperatures. The relationship of these two symmetries, and their transitions around the critical point are important to the continued study of high temperature QCD and deconfined matter which were present at early times in the universe.
@@ 2.2 Emergence of $SU(2)_{CS}$
The set of the generators of the so-called $SU(2)_{CS}$ chiral-spin or $SU(2)_{CS}$ symmetry, explored in [18, 17, 19, 28, 29]. However, for the case of QCD with quarks interacting with the gluonic medium the symmetry does not appear as directly as the free quark example above, and appears to be approximate. This is due to contributions from the chromoelectic term of the quark-gluon interaction(when fixed to the rest frame of the medium), as this term is symmetric under the same transformations $\Sigma$. However, the symmetry breaks down at higher temperatures due the increase in chromomagnetic effects and thus the emergent symmetry of the chromoelectic and quarks field disappears in the high temperature limit[28, 1, 2, 5].
@@ 3 Numerical Results
For fittings which have significant noise, both the effective mass fits as well as correlator fits were used to determine a reliable fitting range.
@@
4 Summary of Preliminary Results
The most significant difference between our work and the results by HotQCD[30], plotted as shaded bands in figs 1 and 2, is the behavior of the $S$ channel at low temperatures. Because our simulations are done with Möbius Domain Wall fermions we do not observe the same $2\pi^{\pm}$ lattice artifact, and instead, we observe a large increase in the screening mass in the region of broken chiral symmetry.
In contrast to both of these symmetries we also plotted a temperature spectrum for the $SU(2)_{CS}$ predicted in [16, 17, 18], unlike the clear signals of symmetry restoration seen in figs 3 and 4, the plot of the symmetry pair $X-A$ figure 4 does not converge to zero for the temperatures in our study. While there is a significant reduction in noise, we do not see a convergence to, or sustained, zero value indicating emergence of $SU(2)_{CS}$.